### v1

# Nandan Kumar Jha

Abstract
LayerNorm is a critical component in modern large language models (LLMs) for stabilizing training and ensuring smooth optimization. However, it introduces significant challenges in mechanistic interpretability, outlier feature suppression, faithful signal propagation, and computational and communication complexity of private inference. This work explores desirable activation functions in normalization-free decoder-only LLMs. Contrary to the conventional preference for the GELU in transformer-based models, our empirical findings demonstrate an opposite trend—ReLU significantly outperforms GELU in LayerNorm-free models, leading to an 8.2% perplexity improvement. We discover a key issue with GELU, where early layers experience entropic overload, leading to the under-utilization of the representational capacity of attention heads. This highlights that smoother activations like GELU are ill-suited for LayerNorm-free architectures, whereas ReLU’s geometrical properties—specialization in input space and intra-class selectivity—lead to improved learning dynamics and better information retention in the absence of LayerNorm. This study offers key insights for optimizing transformer architectures where LayerNorm introduces significant challenges.
1 Introduction
Motivation and challenges. LayerNorm [1] has been a key architectural component contributing to the success of large language models (LLMs) by stabilizing training through normalizing inputs across features within a layer. Additionally, it plays a crucial role in enhancing the models’ non-linear representational capabilities [2, 3, 4, 5]. Despite its benefits, LayerNorm introduces several practical challenges that become pronounced in specific settings:
  1. 1.
    Private Inference (PI): PI protocols enable inference on encrypted data without exposing inputs, ensuring data privacy while protecting model weights [6, 7, 8, 9, 10, 11, 12, 13, 14]. Hybrid PI protocols encounter difficulties with the inverse square root computation inherent in LayerNorm, making it the second most costly operation after GELU, contributing to 22% of total latency and communication overhead [11]. Also, Homomorphic Encryption (HE)-only PI requires polynomial approximations of LayerNorm, which are challenging due to the wide variance range [8].
  2. 2.
    Mechanistic Interpretability: LayerNorm increases the complexity of the residual stream, making it harder to analyze and understand the internal workings of transformer models [15], limiting the applicability of LLMs for applications requiring transparency and explainability.
  3. 3.
    Low-Precision Training: The trainable parameters in LayerNorm are associated with the amplification of outlier features which poses challenges in LLM quantization, as it exacerbates numerical instability and degrades performance in low-precision training regimes [16, 17, 18, 19].
  4. 4.
    Signal Propagation: LayerNorms shown to negatively impact the faithful signal propagation [20].
These challenges highlight the need for LayerNorm-free architectures that preserve transformer benefits while avoiding its drawbacks. However, this shift introduces new considerations, especially in selecting activation functions for the feed-forward networks (FFNs). Prior work [20, 21, 22] has explored various architectural heuristics for designing normalization-free LLMs. However, the impact of removing normalization layers on the choice of FFN activation functions remains underexplored.
(a) SM + LN + G
(b) SM + LN + R
(c) SM + G
(d) SM + R
Figure 1: Entropy heatmaps of attention for baseline (a, b) and normalization-free (c, d) GPT-2 models with GELU and ReLU in the FFN. In the absence of LayerNorm, GELU in the FFN leads to significantly higher entropic overload (highlighted in yellow, c) compared to ReLU.
Research insights and its implications. In this work, we go beyond prior approaches by conducting an in-depth investigation into the design choices for activation functions in normalization-free LLMs, offering new insights into how these choices impact learning dynamics, internal representations, and overall model performance. Our study reveals several key findings:
  • ReLU Outperforms GELU in LayerNorm-Free Models: Contrary to conventional practices, we show that models using ReLU in the FFN significantly outperform, with an 8.2% improvement in perplexity, those using GELU in the absence of LayerNorm (see Figure 3 and Table 3).
  • Learning Dynamics with Learnable Negative Slopes: To explore further, we experimented with a learnable negative slope in the leaky ReLU activation function using two configurations: (1) Layer-wise configuration: Each layer has its independent learnable slope. Initially, early layers learn a positive slope while deeper layers learn a negative slope. However, gradually all layers converge to a near-zero slope (Figure 3a). (2) Global Configuration: A single learnable slope is shared across all layers. The slope initially shifts to positive before converging to near zero (see Figure 3b). These results highlight LayerNorm-free models’ inherent preference for ReLU-like activations with zero negative slopes.
  • Entropic Overload with GELU Activation: To delve deeper, we analyze the head-wise entropy values and find that early layers in normalization-free models with GELU activation experience entropic overload, a significant proportion of attention heads reach near-maximum entropy levels, indicating the under-utilization of the representational capacity of attention heads.
Contributions. Our key contributions are follows:
  1. 1.
    We conduct an in-depth analysis of activation functions in normalization-free, decoder-only models by studying their learning dynamics when trained from scratch.
  2. 2.
    We explore the effect of different activation functions in baseline and normalization-free models on the attention score distribution through the lens of Shannon’s entropy, offering valuable insights for advancing the architectural design of LayerNorm-free models.
  3. 3.
    We conducted experiments across various context sizes (128 and 256) on GPT-2 and Pythia [23] model with 2.1B training tokens from the CodeParrot [24].
2 Preliminaries
Notations. We denote the number of layers as $L$, number of heads as $H$, model dimensionality as $d$, head dimension as $d_{k}$ (where $d_{k}=\frac{d}{H}$), and context length as $T$. Table 1 illustrates the abbreviations for architectural configurations with simplified nonlinearities in a transformer-based LLM.
An overview of transformer-based decoder-only architecture. A transformer-based LLM is constructed by sequentially stacking $L$ transformer blocks, where each block is composed of two sub-blocks: an attention mechanism and a feed-forward network (FFN), both having their own residual connections and normalization layers, positioned in the Pre-LN order to improves training stability [25]. Formally, transformer blocks take an input sequence $\mathbf{X}_{\text{in}}\in\mathbb{R}^{T\times d}$, consisting of $T$ tokens of dimension $d$, and transform it into $\mathbf{X}_{\text{out}}$ as follows:
$\mathbf{X}_{\text{out}}=\hat{\mathbf{X}}_{\text{SA}}+\text{FFN}_{\text{GELU}}(\text{LayerNorm}_{2}(\hat{\mathbf{X}}_{\text{SA}})),\;\text{where}\;\hat{\mathbf{X}}_{\text{SA}}=\mathbf{X}_{\text{in}}+\text{MHA}(\text{LayerNorm}_{1}(\mathbf{X}_{\text{in}})).$ (1)
The Multi-Head Attention (MHA) sub-block enables input contextualization by sharing information between individual tokens. MHA employs the self-attention mechanism to compute the similarity score of each token with respect to all other tokens in the sequence. In particular, self-attention mechanism transform the input sequence $\mathbf{X}$ into $\mathbf{Attn}(\mathbf{X})$ as follows:
$\mathbf{Attn}(\mathbf{X})=\Big{(}\text{Softmax}\Big{(}\frac{1}{\sqrt{d_{k}}}(\mathbf{X}\mathbf{W}^{Q})(\mathbf{X}{\mathbf{W}^{K}})^{\top}+\mathbf{M}\Big{)}\Big{)}\mathbf{X}\mathbf{W}^{V}.$ (2)
Here, each token generates query($Q$), key($K$), and value($V$) vectors through the linear transformations $\mathbf{W}^{Q},\mathbf{W}^{K},\;\text{and}\;\mathbf{W}^{V}\in\mathbb{R}^{d\times d}$, respectively. Then, similarity scores are computed by taking the dot product of the $Q$ and $K$ vectors, scaled by the inverse square root of the $K$ dimension, and passed through a softmax function to obtain the attention weights. These weights are then used to compute a weighted sum of the $V$ vectors, producing the output for each token. For auto-regressive models (e.g., GPT), mask $\mathbf{M}\in\mathbb{R}^{T\times T}$, which has values in $\{0,-\infty\}$ with $\mathbf{M}_{i,j}=0\,\text{iff}\,{i\geq j}$, is deployed to prevent the tokens from obtaining information from future tokens.
The MHA sub-block employs a self-attention mechanism across all the heads, each with its own sets of $Q$, $K$, and $V$. This allows the attention heads to focus on different parts of the input sequence, capturing various aspects of the input data simultaneously. The outputs from all heads are concatenated and linearly transformed ($\mathbf{W}^{O}\in\mathbb{R}^{d\times d}$) to produce the final MHA output as follows:
$\text{MHA}(\mathbf{X})=\text{Concat}\big{(}\text{Attn}_{1}(\mathbf{X}),\;\text{Attn}_{2}(\mathbf{X}),\;\text{Attn}_{3}(\mathbf{X}),\dots,\text{Attn}_{H}(\mathbf{X})\big{)}\mathbf{W}^{O}.$ (3)
Following the MHA sub-block, the FFN sub-block transforms each token independently. The FFN sub-blocks have a single hidden layer whose dimension is a multiple of $d$ (e.g., $4d$ in GPT [26] models). Specifically, the FFN sub-block first applies a linear transformation to the input $\mathbf{X}$ using $\mathbf{W}^{\text{ffn}}_{\text{in}}\in\mathbb{R}^{d\times 4d}$, followed by a non-linear transformation using an activation function such as GELU. This is then followed by another linear transformation using $\mathbf{W}^{\text{ffn}}_{\text{out}}\in\mathbb{R}^{4d\times d}$, as follows:
$\text{FFN}(\mathbf{X})=(\text{GELU}(\mathbf{X}\mathbf{W}^{\text{ffn}}_{\text{in}}))\mathbf{W}^{\text{ffn}}_{\text{out}}$ (4)
2.1 Entropy as a Metric for Attention Score Distribution
Shannon’s entropy quantifies the uncertainty in a probability distribution, measuring the amount of information needed to describe the state of a stochastic system [27, 28]. For a probability distribution $P(x)$, the entropy is defined as $\mathbf{E}(P)=-\sum_{i}P(x_{i})\log P(x_{i})$. Refer to [29] for details on entropy.
In a softmax-based attention mechanism, each softmax operation yields an entropy value representing the sharpness or spread of the attention scores for each query position [30, 31]. Higher entropy indicates a more uniform distribution of softmax scores, while lower entropy signifies a more focused distribution on certain features [32].
Let $\mathbf{A}^{(h,l)}\in\mathbb{R}^{T\times T}$ be the attention matrix of $h$-th head in $l$-th layer, and each element in the attention matrix, $a_{ij}^{(l,h)}$, are attention weights, which are non-negative and sum to one for a query:
$\mathbf{A}^{(l,h)}=\left[a_{ij}^{(l,h)}\right]_{T\times T},\quad\text{where}\quad a_{ij}^{(l,h)}\geq 0\quad\text{and}\quad\sum_{j=1}^{T}a_{ij}^{(l,h)}=1$ (5)
This square matrix is generated by applying the softmax operation over the key length for each query position as follows (i.e., $\mathbf{X}\in\mathbb{R}^{T\times T}$ $\mathbf{X}_{i}\in\mathbb{R}^{1\times T}$ ):
$\mathbf{A}^{(h,l)}(\mathbf{X})=\text{Softmax}\Big{(}\frac{1}{\sqrt{d_{k}}}(\mathbf{X}\mathbf{W}^{Q})(\mathbf{X}{\mathbf{W}^{K}})^{\top}\Big{)},\;\text{where}\quad\text{Softmax}(\mathbf{X}_{i})=\frac{\exp\left(x_{i}\right)}{\sum_{j=1}^{T}\exp\left(x_{j}\right)}$ (6)
Thus, each element $a_{ij}^{(l,h)}$ of the attention matrix can be represented as follows:
$a_{ij}^{(l,h)}=\frac{\exp\left(\frac{1}{\sqrt{d_{k}}}(\mathbf{X}_{i}\mathbf{W}^{Q})(\mathbf{X}_{j}\mathbf{W}^{K})^{\top}\right)}{\sum_{k=1}^{T}\exp\left(\frac{1}{\sqrt{d_{k}}}(\mathbf{X}_{i}\mathbf{W}^{Q})(\mathbf{X}_{k}\mathbf{W}^{K})^{\top}\right).}$ (7)
Following [33], we compute the mean of entropy values across all query positions to obtain a single entropy value for each head. The entropy $\mathbf{E}^{(l,h)}$ for the $h$-th head in the $l$-th layer of an attention matrix is given by:
$\mathbf{E}^{(l,h)}=-\frac{1}{T}\sum_{i=1}^{T}\sum_{j=1}^{T}a_{ij}^{(l,h)}\log\left(a_{ij}^{(l,h)}+\epsilon\right)$ (8)
where $\epsilon$ is a small constant added for numerical stability to prevent taking the log of zero.
Table 1: Architectural configurations of nonlinearities in LLMs, illustrating the combinations of Softmax (SM), LayerNorm (LN), GELU (G), and ReLU (R) functions (see Eq. 1, 2, 3 and 4).
Abbreviation Architectural configuration
SM + LN + G $\mathbf{X}_{\text{out}}=\text{FFN}_{\text{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}GELU}}}(\text{{\color[rgb]{.5,0,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,0,.5}LayerNorm}}_{{\color[rgb]{.5,0,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,0,.5}2}}(\text{MHA}(\text{Attn}_{\text{{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}Softmax}}}(\text{{\color[rgb]{.5,0,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,0,.5}LayerNorm}}_{{\color[rgb]{.5,0,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,0,.5}1}}(\mathbf{X}_{\text{in}})))))$
SM + LN + R $\mathbf{X}_{\text{out}}=\text{FFN}_{\text{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}ReLU}}}(\text{{\color[rgb]{.5,0,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,0,.5}LayerNorm}}_{{\color[rgb]{.5,0,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,0,.5}2}}(\text{MHA}(\text{Attn}_{\text{{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}Softmax}}}(\text{{\color[rgb]{.5,0,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,0,.5}LayerNorm}}_{{\color[rgb]{.5,0,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,0,.5}1}}(\mathbf{X}_{\text{in}})))))$
SM + G $\mathbf{X}_{\text{out}}=\text{FFN}_{\text{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}GELU}}}(\text{MHA}(\text{Attn}_{\text{{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}Softmax}}}(\mathbf{X}_{\text{in}})))$
SM + R $\mathbf{X}_{\text{out}}=\text{FFN}_{\text{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}ReLU}}}(\text{MHA}(\text{Attn}_{\text{{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}Softmax}}}(\mathbf{X}_{\text{in}})))$
The combination of MHA and FFN sub-blocks, along with residual connections and normalization layers, allows transformer models to learn the contextual relationships between tokens effectively.
3 Activation Functions and Their Impact Through Shannon’s Entropy
In this section, we investigate the role of activation functions in baseline and normalization-free decoder-only LLMs. Specifically, we examine the learning dynamics and internal representations of activation functions, using entropy as a metric to highlight key observations and insights.
Well-behaved entropy distribution
Figure 2: Entropy distribution across various configurations of nonlinearities in GPT-2.
We begin by analyzing the headwise entropy distribution of baseline architecture with GELU and ReLU in the FFN, i.e., configurations ${\tt SM+LN+G}$ and ${\tt SM+LN+R}$ respectively. We find that the majority of heads ($\approx$90%) possess entropy values between $\frac{\text{max}}{4}$ and $\frac{\text{3max}}{4}$, where ${\tt max}$ is maximum observed entropy value among all heads (Figure 2). This concentration in the middle entropy range, while avoiding extremes, demonstrates a well-behaved distribution, providing a benchmark for assessing the impact of activation functions on model behavior.
Entropic overload We observed that in certain nonlinearity configurations, a disproportionately large fraction of the attention heads exhibit higher entropy values (between $\frac{3\text{max}}{4}$ and ${\tt max}$). We term this phenomenon as entropic overload and hypothesize that this imbalance results in under-utilization of the network’s representational capacity, as too many heads engaged in exploration, hindering the model from effectively leveraging the diversity of attention heads.
To investigate further, we examined how entropy values evolve during training. Typically, all heads start with higher entropy values, indicating an initial exploration phase, and gradually adapt to balance exploration and exploitation in baseline networks (see Figure 4). However, in the absence of certain nonlinearities, this balance is disrupted, preventing attention heads from specializing and refining their focus on critical aspects of the input, thereby diminishing overall performance.
[…]
figureEvaluation loss for baseline and normalization-free GPT-2 models with GELU and ReLU activations.
Configurations Eval PPL +$\Delta$(%)
SM + LN + G 2.69 0.00
SM + LN + R 2.76 2.53
SM + G 3.20 18.92
SM + R 2.94 9.20
[…]
tablePerplexity comparison of baseline and normalization-free GPT-2 models with GELU and ReLU activations in the FFN, trained from scratch. While the baseline model with GELU outperforms its ReLU counterpart, the normalization-free design shows the opposite trend.
Observation 1: ReLU significantly outperforms GELU in LayerNorm-Free LLMs. While GELU is typically preferred over ReLU in conventional transformer-based models due to its smooth and differentiable properties that improve performance and optimization, our empirical findings indicate the opposite trend for LayerNorm-free models— using ReLU in the FFN exhibit better learning dynamics than their GELU counterpart. This leads to an 8.2% improvement in perplexity for GPT-2 (see Figure 3 and Table 3). A similar trend has been observed on the LN-free Pythia-70M model across various context lengths (see Table 2).
(a) Layerwise learnable slope
(b) Global learnable slope
Figure 3: Learnable negative slope for leaky ReLU in FFN of LN-free GPT-2 model. (a) Layerwise slopes showing initial variability and convergence towards zero. (b) Global slope trend towards zero over training steps, indicating a preference for zero negative slope in LN-free architectures.
To further strengthen our findings, we conducted experiments with a learnable negative slope in the leaky ReLU activation function with two configurations: 1) layer-wise, where each layer has its independent learnable slope, and 2) global, where a single learnable slope is shared across all layers. Results are shown in Figure 3. Interestingly, in the layerwise setting, the early layers initially learn a positive slope while the deeper layers learn a negative slope. However, as training progresses, all layers converge to a near-zero slope. In the global setting, the slope first shifts to positive before converging to near zero. The layerwise entropy dynamics for both cases are plotted in Figure 5.
This highlights the distinct learning dynamics of nonlinearity choices, and a natural preference for zero negative slope, similar to ReLU, in the FFN activation function of the LN-free model.
Observation 2: Early layers in the LayerNorm-Free model with GELU in FFN experience entropic overload. To understand the zero negative slope preference for the FFN activation function in LN-free architecture, we analyzed the headwise entropy values of LN-free models with GELU and ReLU, when trained from scratch, and compared them to their baseline counterparts. Our analysis revealed a significant divergence in the headwise entropy distributions of the LN-free GELU model (see Figure 1). While baseline models with GELU and ReLU exhibit a balanced entropy distribution, by avoiding the extreme values, the LN-free GELU model shows entropic overload in early layers.
(a) SM + LN + G
(b) SM + LN + R
(c) SM + G
(d) SM + R
Figure 4: Evolution of Layerwise entropy when GPT-2 ($L$=12, $H$=12, $d$=768) models with various nonlinearity configurations are trained from scratch on CodeParrot dataset.
Quantitatively, 58% of heads in the LN-free GELU model have entropy values between $\frac{\text{3max}}{4}$ and ${\tt max}$, compared to only 23% in the LN-free ReLU model (Figure 3a). More importantly, very few heads in the latter approach maximum entropy compared to the former (see yellow regions in Figure 1c), indicating more severe entropic overload in the LN-free model with GELU.
These observations align with geometrical properties of ReLUs: it preserve more information about the structure of the raw input, encouraging neurons to specialize in different regions of the input space, leading to a higher intra-class selectivity and specialization [34]. Thus, the lack of LayerNorm makes the geometry and specialization effects of ReLU more beneficial, while GELU’s smoother nonlinearity causes issues in maintaining distinct attention head behaviors.
4 Conclusion
In this paper, we investigated the design of LayerNorm-free decoder-only language models and highlighted the critical role of activation functions in such architectures. Our empirical studies revealed that, contrary to conventional practices, the ReLU activation significantly outperforms, an 8.2% improvement in perplexity, the GELU in normalization-free models. We found that models with learnable negative slopes in leaky ReLU activations naturally converge toward zero negative slopes, effectively resembling ReLU. Additionally, we discovered that LayerNorm-free models with GELU activation suffer from entropic overload in early layers, leading to under-utilization of their representational capacity. These findings underscore the necessity of rethinking activation function choices when LayerNorm is absent and suggest that selecting appropriate activations like ReLU enables the development of transformer models that are more efficient, interpretable, and better suited for applications such as private inference and quantization.
Limitations. This study mainly focuses on pre-training performance, with perplexity as the primary metric, and does not include experiments to evaluate other capabilities such as transfer learning or few-shot learning. Additionally, our findings are been validated on models with fewer than 1B parameters. Future work will explore broader experimental evaluations, including the large-scale models (see Appendix D).
Appendix
[…]
Appendix A Why Use Entropy to Evaluate the Impact of Nonlinearities?
We use entropy as a metric to study the impact of nonlinearities on the transformer-based LLMs for the following reasons:
  • Quantifying attention distribution: As the attention mechanism is fundamental to all transformer-based architecture, computing the entropy of attention score distributions reveals how nonlinearities affect attention concentration. High entropy quantifies exploration and low entropy indicates exploitation.
  • Feature selection: Nonlinearities like ReLU enable feature selectivity by amplifying important features and suppressing less relevant ones [35]. Entropy can measure this selectivity across layers and heads, providing insights into the model’s prioritization of information. Previously, entropy has been used to quantify the layerwise information flow in neural networks [36].
  • Exploration vs. exploitation: Nonlinear operators like the self-attention mechanism, LayerNorm, and GELU balance exploration and exploitation by selecting relevant features while considering a broader context. For instance, heads in the first layer focus on exploration, while those in the second layer focus on exploitation. (see Figures 1a, 1b, 4a and 4b).
  • Systematic assessment: Prior work [37, 32, 33, 31, 30] also used entropy to analyze the behavior of transformer-based models; thus, enhancing validity and comparability of our findings.
Appendix B Design of Experiments
Datasets We train models from scratch using the CodeParrot [24] dataset. The CodeParrot dataset, sourced from 20 million Python files on GitHub, contains 8 GB of files with 16.7 million examples, each with 128 tokens, totaling 2.1B training tokens. We use a tokenizer with a vocabulary of 50K.
Training settings For training on the CodeParrot dataset, we adopt the training settings from [22], which remain consistent across all architectural variations to accurately reflect the impact of the architectural changes. Following [22, 38, 39], all the models are trained on a single RTX 3090 GPU.
B.1 Perplexity as a Reliable Metric to Evaluate the LLMs’ Performance
Perplexity [40] is a widely adopted metric to evaluate the predictive performance of auto-regressive language models, reflecting the model’s ability to predict the next token in a sequence. However, for perplexity to serve as a meaningful comparative metric across different architectures, it is critical to ensure consistency in the tokenizer, and vocabulary size and quality [41]. Any variation in these components can potentially skew the results by inflating or deflating perplexity scores; thus, obfuscating the true effects of architectural changes.
In our work, we maintain tokenization schemes and vocabulary attributes as invariant factors across all experiments within a dataset. This isolation of architectural modifications ensures that any observed variations in perplexity are directly attributable to changes in the model design. Thus, by enforcing a consistent tokenization scheme and vocabulary within a dataset, we ensure that perplexity remains a reliable metric for comparing model architectures. Consequently, lower perplexity in our evaluations reliably reflects improved token-level predictions.
B.2 Why Training from Scratch to Study Nonlinearities?
Understanding the intricate roles of architectural components and nonlinearities—such as activation functions (e.g., GELU, ReLU) in FFN, normalization layers (e.g., LayerNorm), etc.—in transformer-based language models necessitates a methodical and detailed investigative approach. Training models from scratch is essential for this purpose, as it allows us to delve into the internal mechanisms of the model using quantitative measures like entropy. Below, we present a justification for our methodology:
  • Nonlinearities’ impact on the fundamental learning dynamics: Nonlinearities significantly influence the optimization landscape by affecting gradient flow and the model’s ability to navigate non-convex loss surfaces. Training models from scratch allow us to observe the fundamental learning dynamics that emerge during the initial stages of training. Thus, constructing models with controlled variations, such as substituting or excluding specific nonlinearities, enables us to isolate their direct effects impact on convergence behavior and training stability.
  • Understanding internal mechanisms through entropy analysis: Training from scratch enables us to navigate the evolution of entropy values across the layers and assess how architectural components influence information flow within the model. This analysis provides deep insights into the internal workings of models that may not be accessible when starting from pre-trained checkpoints.
  • Limitations of fine-tuning approaches: The aforementioned granular level of analysis is unattainable when starting from pre-trained models, where the optimization trajectory has already been largely determined. In contrast, training models from scratch eliminates confounding variables that could arise from pre-existing weights and learned representations, ensuring that any observed effects are solely due to the architectural modifications introduced.
Appendix C Additional Results
C.1 Layerwise Entropy Dynamics for Learnable Negative Slope
(a) Layerwise learnable slope
(b) Global slope
Figure 5: Evolution of layerwise entropy in LayerNorm-free GPT-2 models ($L$ = 12, $H$ = 12, $d$ = 768) with a learnable negative slope in the leaky ReLU activation function, trained from scratch on the CodeParrot dataset
C.2 Results on Pythia-70M
Table 2: Results for nonlinearity reduction in Pythia-70M ($L$=6, $H$=8, $d$=512) on CodeParrot dataset. LN-free architecture with ReLU activations in FFN (i.e., ${\tt SM+R}$) significantly outperform their GELU counterpart (i.e., ${\tt SM+G}$) across various content length ($T$).
$T$=128 $T$=256
Eval PPL +$\Delta$(%) Eval PPL +$\Delta$(%)
SM+LN+G 3.512 0.00 3.054 0.00
SM+LN+R 3.590 2.22 3.107 1.73
SM+G 4.086 16.35 3.570 16.87
SM+R 3.736 6.36 3.273 7.17
Appendix D Future Work
Scaling up and generalizing to larger models This research opens several avenues for optimizing LayerNorm-free transformer architectures. A primary direction is scaling up experiments to larger models. Evaluating whether the benefits of ReLU activation persist in models with significantly more parameters will determine the applicability of our findings to state-of-the-art language models. Additionally, extending the analysis to other architectures, such as encoder-only or encoder-decoder transformers, could help generalize our insights across different model types.
Longitudinal studies and downstream task performance To validate the long-term viability of this approach, conducting longitudinal studies is crucial. Ensuring that the performance benefits observed with ReLU activation persist over extended training periods and across different tasks would reinforce our findings. Furthermore, while our study focused on perplexity and entropy metrics, future work should analyze how the choice of activation function affects performance on various downstream tasks. Investigating the implications for fine-tuning processes could also provide valuable insights for practical applications.
Hybrid activation strategies Exploring hybrid activation strategies presents another promising avenue. By using different activation functions in different parts of the model—such as combining ReLU in the earlier layers with GELU in the later layers—we could strike a balance between the benefits observed in our study and the traditional advantages of GELU. This approach may enhance model performance while maintaining computational efficiency.
Interpretability and practical applications Given the observed differences in entropy distribution between ReLU and GELU models, future research could explore how different activation functions impact the interpretability of LayerNorm-free models. This could potentially lead to more explainable AI systems, addressing a critical need in the field. Integrating these findings into practical applications like private inference and quantization is also promising, as it could improve both model efficiency and security.
Knowledge distillation for performance enhancement By distilling knowledge from a larger, LayerNorm-equipped teacher model to a smaller, LayerNorm-free student model with appropriate activation functions, the performance and generalization capabilities of the student model can be improved. This approach could mitigate any performance gaps arising from the absence of LayerNorm while maintaining the benefits of simplified computation and improved interpretability.